OpenStudy (anonymous):

How do you solve this using substitution? x squared - 2x + y squared 2x + y = 4

OpenStudy (anonymous):

Easy (by rearranging) x^2 - y^2 = 2x

OpenStudy (anonymous):

U following so far?

OpenStudy (anonymous):

So if 2x=x^2 -y^2 thhheeeeennnnn X squared minus y squared plus y=4

OpenStudy (anonymous):

Oh wait, I forgot something. It should x squared -2x + y squared = 3

OpenStudy (anonymous):

Here this will be a lot simpler If 2x + y = 4 then y =4-2x

OpenStudy (anonymous):

Now since y =4-2x then by substituting y into the other equation of x squared minus 2x plus y squared = 3

OpenStudy (anonymous):

After substitution you're answer new equation will be x^2 -2x + (4-2x)^2=3

OpenStudy (anonymous):

Now solve for x

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

Do I need to subtract 3 and put it on the left side to make the equation equal to 0?

OpenStudy (anonymous):

You want to try to get x so yea

OpenStudy (anonymous):

So now I should have 5x squared - 18x + 13 = 0 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

And now I need to factor that equation out ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So now I have (5x - 13) (x - 1) which makes x = 13\5 and 1

OpenStudy (anonymous):

So that's youre x and you just substitute those into the equation for y which is y=4-2x

OpenStudy (anonymous):

y = 2 and -6\5

OpenStudy (anonymous):

There you go those are your answers

OpenStudy (anonymous):

You get two answers because technically you have a parabola and a line and your asking where they intersect and they intersect at two points

OpenStudy (anonymous):

Do u get that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Glad to help then :)